# pMixFed: Efficient Personalized Federated Learning through Adaptive Layer-Wise Mixup

### v1

@@ 10.1 Experimental Setup
Dataset: We used three widely used federated learning datasets: MNIST [23], CIFAR-10, CIFAR-100 [3], and EMNIST [8]. CIFAR-10 consists of 50,000 images of size $32\times 32$ for training and 10,000 images for testing. CIFAR-100, on the other hand, consists of 100 classes, with 500 $32\times 32$ images per class for training and 100 images per class for testing. MNIST contains 10 labels and includes 60,000 samples of $28\times 28$ grayscale images for training and 10,000 for testing. For creating heterogenity, we followed the
@@
Training Details: For evaluation, We have reported the average test accuracy of the global model [59] for different approaches. The final global model at the last communication round is saved and used during the evaluation. The global model is then personalized according to each baseline’s personalization or fine-tuning algorithm for $r=4$ local epochs and $T=50$. For FedAlt, the local model is reconstructed from the global model and fine-tuned on the test data. For FedSim, both the global and local models are fine-tuned partially but simultaneously. In the case of FedBABU, the head (fully connected layers) remains frozen during local training, while the body is updated. Since we could not directly apply pFedHN in our platform setting, we adapted their method using the same hyper parameters discussed above and employed hidden layers with 100 units for the hypernetwork and 16 kernels. The local update process for LG-FedAvg, FedAvg, and Per-FedAvg simply involves updating all layers jointly during the fine-tuning process. The global learning rate for the methods that need sgd update in the global server e.g., FedAvg, has been set from $lr_{global}=[1e-3,1e-4,and1e-5]$.It should be noted that due to the performance drop for some methods (FedAlt , FedSim) in round 10 or 40 in some settings, we’ve reported the highest accuracy achieved. Also this is the reason the accuracy curves are illustrated for 39 rounds instead of 50.  [^5] As discussed in the main paper, the change of hyper-parameters such as lr, batch size, momentum and even changing the optimizer to adam didn’t help with the performance drop in most cases.
@@ 10.2 Out-of-Sample and Participation Gap
In the evaluation of the effect of learning rate and mixup, the average test accuracy [^6] Classification accuracy using softmax is measured on cold-start clients $|D^{ts}_{k\cap\text{unseen}}|,k\subset\{1,\dots,M\}$, where $D^{ts}\neq D^{tr}$. These clients have not participated in the federation at any point during training. FedAlt and FedSim perform poorly on cold-start users or unseen clients, highlighting their limited generalization capability. The test accuracy of pFedMix, while affected under a 10% participation rate, benefits significantly from seeing more clients. Increased client participation directly improves accuracy, as observed in previous studies [34].
@@ 10.3 Ablation Study : The Effects of Different alpha on Mixup Degree
    $\alpha=\begin{cases}0.1,&\text{initial stage (early training)}\\ 100,&\text{middle stage (convergence phase)}\\ 10,&\text{final stage (fine-tuning)}\end{cases}$ (29)
@@
Table 3: Accuracy of $pMixFed$ with Different $\alpha$ Values in the Beta Distribution
@@
Table 4: Accuracy of $pMixFed$ with Different $\alpha$ Values and Sampling Strategies

### v2

@@ 10.1 Experimental Setup
For creating heterogenity, we followed the Training Details: For evaluation, We have reported the average test accuracy of the global model yuan2021we for different approaches. The final global model at the last communication round is saved and used during the evaluation. The global model is then personalized according to each baseline’s personalization or fine-tuning algorithm for $r=4$ local epochs and $T=50$. For FedAlt, the local model is reconstructed from the global model and fine-tuned on the test data. For FedSim, both the global and local models are fine-tuned partially but simultaneously. In the case of FedBABU, the head (fully connected layers) remains frozen during local training, while the body is updated. Since we could not directly apply pFedHN in our platform setting, we adapted their method using the same hyper parameters discussed above and employed hidden layers with 100 units for the hypernetwork and 16 kernels. The local update process for LG-FedAvg, FedAvg, and Per-FedAvg simply involves updating all layers jointly during the fine-tuning process. The global learning rate for the methods that need sgd update in the global server e.g., FedAvg, has been set from $lr_{global}=[1e-3,1e-4,and1e-5]$.It should be noted that due to the performance drop for some methods (FedAlt , FedSim) in round 10 or 40 in some settings, we’ve reported the highest accuracy achieved. Also this is the reason the accuracy curves are illustrated for 39 rounds instead of 50.  [^7] As discussed in the main paper, the change of hyper-parameters such as lr, batch size, momentum and even changing the optimizer to adam didn’t help with the performance drop in most cases.
10.2 Caltech-101 Experimental Setup and Results
Dataset.
Caltech-101 fei2004learning contains 9,146 images across 101 object categories plus a background class, with $\sim$40–800 images per class (most near 50). Images are resized to $224\times 224$ for CNN training. We followed prior FL literature fallah2020personalized; yang2023fedrep; pillutla2022federated in simulating heterogeneous splits: each client receives at most $S=30$ classes sampled by Dirichlet distribution ($\alpha=0.5$). We evaluate four federated settings: $N=\{10,100\}$ clients with participation $C=\{10\%,100\%\}$.
Model and Training.
For consistency with CIFAR/MNIST, we adopt a CNN backbone (4 convolutional layers + 1 fully connected layer). Each client trains for $r=4$ local epochs per round with batch size 32. The total communication rounds are $T=50$, using Adam optimizer with learning rate $1\times 10^{-3}$. For partial baselines (FedAlt), the split layer is fixed at the midpoint. For Per-FedAvg, meta-initialization is updated with inner learning rate $0.01$ and outer learning rate $0.001$. FedRep uses a shared representation with personalized heads (dimension 256). Evaluation is done by fine-tuning on local test splits for 4 epochs.
Results.
Table 6 reports detailed top-1 accuracy (%) across four evaluation angles (A–D), along with the average. Across all federated settings, pMixFed achieves the best performance, with particularly strong improvements in the heterogeneous regime ($N=100$, $C=10\%$). FedRep performs second-best overall, confirming the utility of representation sharing, while Per-FedAvg is strong when client counts are small. FedAlt trails consistently, highlighting the limitations of fixed partitioning.
10.3 Results on Caltech-101 dataset
We additionally evaluate on Caltech-101 fei2004learning, following the same heterogeneous FL setup as CIFAR/MNIST. Each client receives at most $S=30$ classes, and data is split using a Dirichlet distribution ($\alpha=0.5$). We report accuracy across four evaluation angles (A–D) and their mean (AVG).
Table 6: Performance (%) on Caltech-101 (CNN backbone). Rows A–D denote evaluation angles; AVG is the mean across A–D.
$N=100,\penalty 10000\ C=100\%$
Method A B C D AVG
pMixFed (Ours) 78.4 80.2 79.1 79.5 79.3
FedRep 75.8 78.5 77.2 76.6 77.0
Per-FedAvg 75.1 77.2 76.0 77.5 76.5
FedAlt 70.8 74.1 73.5 73.9 73.1
$N=10,\penalty 10000\ C=100\%$
Method A B C D AVG
pMixFed (Ours) 80.9 83.8 82.9 81.7 82.4
FedRep 78.6 80.9 80.4 80.5 80.1
Per-FedAvg 79.0 81.8 82.1 81.0 81.0
FedAlt 74.0 76.8 75.2 76.0 75.5
$N=100,\penalty 10000\ C=10\%$
Method A B C D AVG
pMixFed (Ours) 74.6 77.8 76.9 74.9 76.1
FedRep 72.9 74.6 74.5 74.8 74.2
Per-FedAvg 72.1 74.0 73.8 74.1 73.5
FedAlt 67.4 70.8 69.9 69.8 69.5
$N=10,\penalty 10000\ C=10\%$
Method A B C D AVG
pMixFed (Ours) 77.2 80.1 79.8 79.0 79.0
FedRep 75.1 77.3 77.2 77.8 76.9
Per-FedAvg 74.2 76.5 76.0 77.9 76.2
FedAlt 70.5 72.6 72.9 72.8 72.2
@@ 10.4 Out-of-Sample and Participation Gap
In the evaluation of the effect of learning rate and mixup, the average test accuracy [^8] Classification accuracy using softmax is measured on cold-start clients $|D^{ts}_{k\cap\text{unseen}}|,k\subset\{1,\dots,M\}$, where $D^{ts}\neq D^{tr}$. These clients have not participated in the federation at any point during training. FedAlt and FedSim perform poorly on cold-start users or unseen clients, highlighting their limited generalization capability. The test accuracy of pFedMix, while affected under a 10% participation rate, benefits significantly from seeing more clients. Increased client participation directly improves accuracy, as observed in previous studies pillutla2022federated.
@@ 10.5 Ablation Study : The Effects of Different alpha on Mixup Degree
[…]
@@
Table 7: Accuracy of pMixFedpMixFed with Different $\alpha$ Values in the Beta Distribution
@@
Table 8: Accuracy of pMixFedpMixFed with Different $\alpha$ Values and Sampling Strategies